Consider 3 distinct points M1, M2, and M3 with respective affixes z1, z2, and z3. Show that (Triangle M1, M2, and M3 is equilateral) ↔ ( z12 + z22 +z32 - z1z2 - z2z3 - z1z3).




1
Expert's answer
2021-09-27T15:55:05-0400

The problem is similar to this problem:

Given three complex numbers "z_{1}, z_{2}, z_{3}" prove that the points "z_{1}, z_{2}, z_{3}" are vertices of an equilateral triangle in "\mathbb{C}" , if "z_{1}^{2}+z_{2}^{2}+z_{3}^{2}=z_{1} z_{2}+z_{1} z_{3}+z_{2} z_{3}"  .


Three points "z_{1}, z_{2}, z_{3}" are vertices of an equilateral triangle

 

"\begin{aligned}

&\Longleftrightarrow \frac{z_{3}-z_{1}}{z_{2}-z_{1}}=\cos \left(\pm 60^{\circ}\right)+i \sin \left(\pm 60^{\circ}\right)=\frac{1 \pm \sqrt{3} i}{2} \\

&\Longleftrightarrow \quad 2 z_{3}-z_{1}-z_{2}=\pm \sqrt{3} i\left(z_{2}-z_{1}\right) \\

&\Longleftrightarrow\left(2 z_{3}-z_{1}-z_{2}\right)^{2}=\left(\pm \sqrt{3} i\left(z_{2}-z_{1}\right)\right)^{2} \\

&\Longleftrightarrow \quad z_{1}^{2}+z_{2}^{2}+z_{3}^{2}=z_{1} z_{2}+z_{2} z_3+z_{3} z_{1}

\end{aligned}"